OpenStudy (anonymous):

subtract! 3/-5 - 9/4

sammixboo (sammixboo):

Yeah I can help you on this. Hold on a second ppllleeaassee >:)

OpenStudy (anonymous):

no problem...i love how you keep putting smileys XD

OpenStudy (anonymous):

\[\frac{ 3 }{ -5 }-\frac{ 9 }{ 4 }\]

OpenStudy (anonymous):

thats what it looks like

OpenStudy (anonymous):

@sammixboo r you there??

sammixboo (sammixboo):

BACK omg my internet went down. I was just fixing to post the problem step by step, but my internet went bye bye lmao. But what I got was -2 17/20 \(\sf\large(\)\(\large\sf\color{magenta}{-2\frac{17}{20}}\)\(\sf\large)\)

OpenStudy (anonymous):

thats the answer??

sammixboo (sammixboo):

Yeppers :D

OpenStudy (anonymous):

okkk lemmie check..dont move!

OpenStudy (anonymous):

its right! :D can i ask one more??

sammixboo (sammixboo):

See you can count all the negative signs. Once you are done counting (Don't include the subtracting sign) and if the number of negative signs is odd, then the answer will be negative. If the number of negative signs is even, then your answer is positive. Easy tip I learned that helped.

sammixboo (sammixboo):

And yeah. Go ahead.

OpenStudy (anonymous):

\[\frac{ 2 }{ -5}+\frac{ -3 }{ 8 }\]

sammixboo (sammixboo):

Will the answer be positive or negative? (Use the hint I gave you above)

OpenStudy (anonymous):

it will beee...i think positive since two negatives make a positive

sammixboo (sammixboo):

Hold on a second this problem is different with that rule

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

are you alive?? lmao

sammixboo (sammixboo):

Okay I figured it out. I just confused myself. The hint I gave you does work with this lmao. Sorry for the confusion. I was doing the problem wrong. But here is my steps: \(\sf\huge\color{magenta}{\frac{2}{-5}}\) \(\large -\) \(\sf\huge\color{magenta}{\frac{-3}{8}}\) The common denominator is 40 so we must multiply \(\sf\large\color{magenta}{\frac{2}{-5}}\) by 8, and multiply \(\sf\large \color{magenta}{\frac{-3}{8}}\) by 5 \(\sf\huge\color{magenta}{\frac{2x8}{-5x8}}\) \(\large -\) \(\sf\huge\color{magenta}{\frac{-3x5}{8x5}}\) \(\sf\huge\color{magenta}{\frac{-16}{40}}\) \(\large -\) \(\sf\huge\color{magenta}{\frac{-15}{40}}\) \(\sf\huge\color{magenta}{\frac{-31}{40}}\) (SORRY IF THIS IS TAKING FOREVER TO LOAD)

sammixboo (sammixboo):

The subtraction sign is supposed to be an addition sign o.o Sorry I just woke up

OpenStudy (anonymous):

can that be simplified any further and its ok no judgement here

sammixboo (sammixboo):

Nope. It cannot be simplified any further

OpenStudy (anonymous):

ok lemmie check it :)

sammixboo (sammixboo):

Because only 1 and -31 can go into -31

sammixboo (sammixboo):

Okay tyt